OpenStudy (anonymous):

If U = 5i - 5j and V = =-4j + 3j, find W = 5u - 2v

OpenStudy (anonymous):

i have no idea what formula to use

OpenStudy (amistre64):

what is u times 5? or scaled by 5?

OpenStudy (amistre64):

what is v scaled by -2?

OpenStudy (hba):

Vectors :)

OpenStudy (amistre64):

u = <5,5> v = <-4,3> find W = 5u - 2v 5u = 5<5,5> -2v = -2<-4,3> 5u = <25,25> -2v = < 8, -6> ---------------- w = < ?, ? >

OpenStudy (anonymous):

So im dealing with vectors, so do I multiply the i and J's together?

OpenStudy (amistre64):

the i and j are just notation for place holder; you can convert it back to a more usual vector notation as i did if it looks better

OpenStudy (amistre64):

u = 5i - 5j v =-4j + 3j w = 5u - 2v 5u = 5( 5i - 5j) +(-2v) =-2(-4j + 3j) 5u = 25i - 25j +(-2v) = 8j - 6j ----------------- w = ? i + ? j

OpenStudy (amistre64):

i see i dropped a negative the first go around :)

OpenStudy (anonymous):

so w = 33i and -31j?

OpenStudy (amistre64):

yes w = 33i - 31j